OpenStudy (anonymous):

solve the system of equations: -2x-y=-3 y=-3x

OpenStudy (anonymous):

y = -3x so -2x - (-3x) = -3 -2x + 3x = -3 x = -3 Sub this into the first equation and you get y = -3 (-3) = 9 Overall, x = -3 and y = 9